OpenStudy (anonymous):

I must compare fractions by converting them into equivalent 5/9 and 9/12

OpenStudy (anonymous):

20/36 vs 27/36 so 9/12 is greater

OpenStudy (anonymous):

5/9= 20/36 9/12= 27/36 the latter is greater :)

OpenStudy (anonymous):

thank you very much